# Regularized KL-Divergence for Well-Defined Function-Space Variational Inference in Bayesian Neural Networks

### v2

@@ 2.2 Regularized KL divergence for Gaussian measures
Our solution to the negative infinite function-space ELBO builds on a regularized KL divergence, which is expressed in terms of Gaussian measures for the variational posterior and prior. We obtain these Gaussian measures from corresponding Gaussian processes (GP). We first discuss under which conditions a GP induces a Gaussian measure, and then present the regularized KL divergence.
Gaussian measures and Gaussian processes
@@
Definition 2.1 (Gaussian measure, Kerrigan et al., (2023), Definition 1).
Let $(\Omega,\mathcal{B},\mathbb{P})$ be a probability space. A measurable function $F:\Omega\mapsto L^{2}(\mathcal{X},\rho)$ is called a Gaussian random element (GRE) if for any $g\in L^{2}(\mathcal{X},\rho)$ the random variable $\langle g,F\rangle$ has a Gaussian distribution on . For every GRE $F$, there exists a unique mean element $m\in L^{2}(\mathcal{X},\rho)$ and a finite trace linear covariance operator $C:L^{2}(\mathcal{X},\rho)\mapsto L^{2}(\mathcal{X},\rho)$ such that $\langle g,F\rangle\sim\mathcal{N}\mathopen{}\left(\langle g,m\rangle,\langle Cg,g\rangle\right)\mathclose{}$ for all $g\in L^{2}(\mathcal{X},\rho)$. The pushforward of $\mathbb{P}$ along $F$ denoted $\mathbb{P}^{F}:=F_{\#}\mathbb{P}$ is a Gaussian (probability) measure on $L^{2}(\mathcal{X},\rho)$.
Gaussian measures generalize Gaussian distributions to infinite-dimensional spaces where measures do not have associated densities. Following Wild et al., (2022), we notate the Gaussian measure obtained from the GRE $F$ with mean element $m$ and covariance operator $C$ as $\mathbb{P}^{F}:=\mathcal{N}\mathopen{}\left(m,C\right)\mathclose{}$.
@@
Definition 2.2 (Regularized KL divergence, Quang, (2022) Definition 5).
Let $\nu_{1}=\mathcal{N}\mathopen{}\left(m_{1},C_{1}\right)\mathclose{}$ and $\nu_{2}=\mathcal{N}\mathopen{}\left(m_{2},C_{2}\right)\mathclose{}$ be two Gaussian measures with $m_{1},m_{2}\in L^{2}(\mathcal{X},\rho)$ and $C_{1},C_{2}$ bounded, self-adjoint, positive and trace-class linear operators on $L^{2}(\mathcal{X},\rho)$. Let $\gamma\in$, $\gamma>0$ be fixed. The regularized KL divergence is defined as follows,
$D^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\nu_{1}\,\|\,\nu_{2}\right)\mathclose{}:=$ $\,\frac{1}{2}\langle m_{1}-m_{2},(C_{2}+\gamma\mathbb{I})^{-1}\mathopen{}\left(m_{1}-m_{2}\right)\mathclose{}\rangle+\frac{1}{2}\operatorname{Tr}_{X}\left[\mathopen{}\left(C_{2}+\gamma\mathbb{I}\right)\mathclose{}^{-1}\mathopen{}\left(C_{1}+\gamma\mathbb{I}\right)\mathclose{}-\mathbb{I}\right]$
$-\frac{1}{2}\log\operatorname{det}_{X}\left[\mathopen{}\left(C_{2}+\gamma\mathbb{I}\right)\mathclose{}^{-1}\mathopen{}\left(C_{1}+\gamma\mathbb{I}\right)\mathclose{}\right].$ (2.8)
For any $\gamma>0$, the regularized KL divergence is well-defined and finite, even if the Gaussian measures are singular (Quang,, 2019). It converges to the conventional KL divergence (if it is well-defined) for $\gamma\to 0$ (Quang,, 2022, Theorem 6). Furthermore, and importantly for our application to VI, if two Gaussian measures $\nu_{1}$ and $\nu_{2}$ are induced by GPs $f_{i}\sim\mathcal{GP}(\mu_{i},K_{i})$ for $i=1,2$, then $D^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\nu_{1}\,\|\,\nu_{2}\right)\mathclose{}$ can be consistently estimated (Quang,, 2022) from a finite number $M$ of samples using the estimator
$\hat{D}^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\nu_{1}\,\|\,\nu_{2}\right)\mathclose{}:=$ $\,\frac{1}{2}\mathopen{}\left({\bm{m}}_{1}-{\bm{m}}_{2}\right)\mathclose{}^{\top}({\bm{\Sigma}}^{(\gamma)}_{2})^{-1}\mathopen{}\left({\bm{m}}_{1}-{\bm{m}}_{2}\right)\mathclose{}+\frac{1}{2}\operatorname{Tr}\big{[}({\bm{\Sigma}}_{2}^{(\gamma)})^{-1}{\bm{\Sigma}}^{(\gamma)}_{1}-\mathbb{I}_{M}\big{]}$
$-\frac{1}{2}\log\det\big{[}({\bm{\Sigma}}_{2}^{(\gamma)})^{-1}{\bm{\Sigma}}^{(\gamma)}_{1}\big{]}$ (2.9)
@@ 3 Generalized function-space VI with the regularized KL divergence
This section presents our proposed generalized function-space variational inference (GFSVI) method, which addresses the problem of the infinite KL divergence discussed in section 2.1, which we take for an indication that VI is too restrictive if one wants to use genuine function-space priors. We instead consider generalized variational inference (Knoblauch et al.,, 2019), which reinterprets the ELBO in Eq. 2.1 as a regularized expected log-likelihood and explores alternative divergences for the regularizer. Specifically, we propose to use the regularized KL divergence. The presentation in this section builds heavily on tools introduced in section 2, which turn out to fit together perfectly: the pushforward of a Gaussian variational distribution in weight-space through the linearized neural network (Eq. 2.6) induces a GP variational posterior (Eq. 2.7) that admits a Gaussian measure on $L^{2}(\mathcal{X},\rho)$ (definition 2.1). Further selecting a GP prior which has an associated Gaussian measure on $L^{2}(\mathcal{X},\rho)$ allows us to use the regularized KL divergence (Eq. 2.8). We present the GFSVI method in section 3.1 and discuss connections to prior work in section 3.2.
3.1 Generalized function-space variational inference
We present a simple and well-defined objective for function-space inference, and an algorithm for its optimization.
Objective function
We start from the ELBO in Eq. 2.3, where we use the Gaussian variational measure $\mathbb{Q}_{\phi}^{F}$ induced by the pushforward of a Gaussian variational distribution $q_{\phi}({\bm{w}})=\mathcal{N}\mathopen{}\left({\bm{w}}\,|\,{\bm{m}},{\bm{S}}\right)\mathclose{}$ along the linearized network $f_{L}$ (Eq. 2.6). The function-space prior may be any GP that has an associated Gaussian measure $\mathbb{P}^{F}$ on $L^{2}(\mathcal{X},\rho)$ (see definition 2.1 and discussion below it). We now replace the KL divergence in the ELBO with the regularized KL divergence $D_{\text{KL}}^{\gamma}$ (Eq. 2.8), which is well-defined and finite for any pair of Gaussian measures. For a likelihood $p\mathopen{}\left(\mathcal{D}\,|\,{\bm{w}}\right)\mathclose{}=\prod_{i=1}^{N}p\mathopen{}\left(y_{i}\,|\,f_{L}({\bm{x}}_{i};{\bm{w}})\right)\mathclose{}$, we thus obtain
@@ 3.1 Generalized function-space variational inference
Input: linearized BNN $f_{L}$ with measure $\mathbb{Q}_{\phi}^{F}$, Gaussian process prior $\mathcal{GP}(\mu,K)$ with measure $\mathbb{P}^{F}$, measurement point distribution $\rho({\bm{x}})$, data $\mathcal{D}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}$, batch size $B$, $\gamma>0$.
for each minibatch $({\bm{x}}_{\mathcal{B}},y_{\mathcal{B}})\sim\mathcal{D}$ do
       Calculate ^$\hat{\ell}_{1}=\frac{N}{B}{\mathop{{}\mathbb{E}}}_{q_{\phi}({\bm{w}})}\mathopen{}\left[\log p\mathopen{}\left(y_{\mathcal{B}}\,|\,f_{L}({\bm{x}}_{\mathcal{B}},{\bm{w}})\right)\mathclose{}\right]\mathclose{}$;
      [-1.5ex] Draw measurement points ${\bm{x}}=\{{\bm{x}}^{(i)}\}_{i=1}^{M}\stackrel{{\scriptstyle\!\!\text{i.i.d.}\!\!}}{{\sim}}\rho({\bm{x}})$;
      [-1.5ex] Calculate ^$\hat{\ell}_{2}=\hat{D}_{\text{KL}}^{\gamma}\big{(}\mathbb{Q}_{\phi}^{F}\,\big{|}\!\big{|}\,{\mathbb{P}^{F}}\big{)}$ using ${\bm{x}}$ (Eq. 2.9);
      [-1.5ex] Calculate ^$\hat{\mathcal{L}}(\phi)=\hat{\ell}_{1}-\hat{\ell}_{2}$;
       Update $\phi$ using a step in the direction $\nabla_{\!\phi}\hat{\mathcal{L}}(\phi)$;
end for
@@
Estimation and optimization
@@
Technical details ($\gamma$ and $\rho$)
It turns out that increasing $\gamma$ reduces the influence of the prior on inference (see figure 17 in Appendix). At the same time, $\gamma$ acts as jitter that prevents numerical errors (see section 3.2). We recommend setting $\gamma$ large enough to avoid numerical errors but sufficiently small to strongly regularize the objective in Eq. 3.1 (see figure 15 in Appendix). We found that GFSVI is robust to a wide range of values and fixed $\gamma=10^{-10}$. Furthermore, the probability measure $\rho$ defined with $L^{2}(\mathcal{X},\rho)$ has to assign non-zero probability to any open set of $\mathcal{X}$ to regularize the BNN on all of its support. Following Rudner et al., 2022b , we draw measurement points from a uniform distribution over $\mathcal{X}$ when using tabular data and explore different configurations (such as samples from other data sets) when using high-dimensional image data (see section A.3.4).
@@ 3.2 Connections to prior work
TFSVI (Rudner et al., 2022b, ) and FVI (Sun et al.,, 2019) solve stability issues by introducing jitter/noise, which has a similar effect as the regularization in Eq. 2.8. However, as mentioned in Section 2.1, TFSVI introduces jitter only to overcome numerical issues and is fundamentally restricted to prior specification in weight space since its function-space prior is the pushforward of a weight-space prior. FVI does not linearize the BNN, and therefore does not have access to an explicit variational measure in function space. This severely complicates the estimation of (gradients of) the KL divergence, and the authors resort to implicit score function estimators, which make their method difficult to use in practice (Ma and Hernández-Lobato,, 2021). Our proposed GFSVI does not suffer from these difficulties as the variational posterior is an explicit (Gaussian) measure. This allows us to estimate the (regularized) KL divergence without sampling any noise or using implicit score function estimators.
@@ 4 Experiments
Figure 2: Inference on synthetic data (gray circles) using an RBF prior for function-space methods GFSVI and FVI. The proposed GFSVI provides the best approximation of the exact GP posterior.
@@
In this section, we evaluate our generalized function-space variational inference (GFSVI) method qualitatively on synthetic data (Section 4.1) and quantitatively on real-world data (Section 4.2). We find that GFSVI accurately captures structural properties specified by the GP prior, that it approximates the exact Gaussian process (GP) posterior more faithfully than BNN baselines, and that it performs competitively on regression, classification and out-of-distribution detection tasks. We also discuss the influence of the BNN’s inductive biases.
Baselines
@@
We compare the proposed GFSVI method to two weight-space inference methods (mean-field variational inference (Blundell et al.,, 2015) and linearized Laplace (Immer et al.,, 2021)) and to three function-space inference methods (FVI (Sun et al.,, 2019), TFVSI (Rudner et al., 2022b, ) and VIP (Ma et al.,, 2019), where TFSVI performs inference in function space but with the pushforward of a weight-space prior and VIP using a BNN prior). All BNNs have the same architecture and fully-factorized Gaussian approximate posterior. We also include results for a sparse GP with a posterior mean parameterized by a neural network (GWI) (Wild et al.,, 2022), as well as for a Gaussian Process (GP) (Williams and Rasmussen,, 2006) (when the size of the dataset allows it), and for a sparse GP (Hensman et al.,, 2013) for regression tasks. Ẇe consider the GP, sparse GP and GWI as gold standards as they represent the exact (or near exact) posterior for models with GP priors.
4.1 Qualitative results on synthetic data
Regression
@@ 4.1 Qualitative results on synthetic data
We consider synthetic regression data with a one-dimensional feature-space $\mathcal{X}\subset$, where the values $y_{i}$ are randomly sampled around $\sin(2\pi x_{i})$ (circles in Figures 1-11, see Appendix A.3.1). The green lines in Figures 1-8 and 11 show functions sampled from the (approximate) posteriors, and the red lines are the inferred mean functions. Figure 2 compares GFSVI with an RBF GP prior to baselines and to the exact posterior. We find that GFSVI visually matches the true posterior best (the same holds for a Matern-1/2 prior, see Figure 5 in the Appendix). Figure 1 in the Introduction and Figure 6 in the Appendix show that GFSVI notably adapts to varying prior assumptions (varying smoothness and length scale, respectively). In addition, Figures 4 and 8 in the Appendix show that GFSVI provides strong regularization when the data generative process is noisy, and that it can be trained with fewer measurement points $M$ than FVI without significant degradation.
Classification
We further consider a binary classification task derived from the 2-dimensional two moons data set (Pedregosa et al.,, 2011), see red and blue dots in Figures 9 and 10 in the Appendix. The first and second row in both figures show the inferred mean probability of class 1 (blue dots) and its 2-standard deviations with respect to posterior samples. Just like for regression, we find that GFSVI captures the beliefs of the RBF and Matern-1/2 GP priors better than all BNN-baselines, and shows greater uncertainty outside of the support of the data than the BNN baselines.
Inductive biases
@@
figure 11 in the Appendix compares GFSVI to the exact GP-posterior across two different priors and three model architectures (details in Appendix A.3.1). We find that, with ReLU activations, small models are prone to underfitting for smooth priors (RBF), and to collapsing uncertainty for rough priors (Matern-1/2). By contrast, with smooth activations (Tanh), smaller models suffice, and they are compatible with most standard GP priors (the results shown in figure 11 extend to RBF, Matern, and Rational Quadratic in our experiments). We also analyzed how the number $M$ of measurement points affects performance. figures 7 and 14 in Appendix show that capturing the beliefs of rough GP priors and estimating $D_{\text{KL}}^{\gamma}$ with these priors requires larger $M$.

### v3

@@ 2.2 Regularized KL divergence
Our solution to the negative infinite function-space ELBO builds on a regularized KL divergence, which is expressed in terms of Gaussian measures for the variational posterior and prior. We obtain these Gaussian measures from GPs. We first discuss under which conditions a GP induces a Gaussian measure, and then present the regularized KL divergence.
Gaussian measures and Gaussian processes.
@@
Definition 2.1 (Gaussian measure, Kerrigan et al. [2023], Definition 1).
Let $(\Omega,\mathcal{B},\mathbb{P})$ be a probability space. A measurable function $F:\Omega\mapsto L^{2}(\mathcal{X},\rho)$ is called a Gaussian random element (GRE) if for any $g\in L^{2}(\mathcal{X},\rho)$ the random variable $\langle g,F\rangle$ has a Gaussian distribution on . For every GRE $F$, there exists a unique mean element $m\in L^{2}(\mathcal{X},\rho)$ and a finite trace linear covariance operator $C:L^{2}(\mathcal{X},\rho)\mapsto L^{2}(\mathcal{X},\rho)$ such that $\langle g,F\rangle\sim\mathcal{N}\mathopen{}\left(\langle g,m\rangle,\langle Cg,g\rangle\right)\mathclose{}$ for all $g\in L^{2}(\mathcal{X},\rho)$. The pushforward of $\mathbb{P}$ along $F$, denoted $\mathbb{P}^{F}:=F_{\#}\mathbb{P}$, is a Gaussian measure on $L^{2}(\mathcal{X},\rho)$.
@@
Definition 2.2 (Regularized KL divergence, Quang [2022] Definition 5).
Let $\nu_{1}=\mathcal{N}\mathopen{}\left(m_{1},C_{1}\right)\mathclose{}$ and $\nu_{2}=\mathcal{N}\mathopen{}\left(m_{2},C_{2}\right)\mathclose{}$ be two Gaussian measures with $m_{1},m_{2}\in L^{2}(\mathcal{X},\rho)$ and $C_{1},C_{2}$ bounded, self-adjoint, positive and trace-class linear operators on $L^{2}(\mathcal{X},\rho)$. Let $\gamma\in{}_{>0}$ be fixed. The regularized KL divergence is defined as follows,
$D^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\nu_{1}\,\|\,\nu_{2}\right)\mathclose{}:=\frac{1}{2}\langle m_{1}-m_{2},(C_{2}+\gamma\mathbb{I})^{-1}\mathopen{}\left(m_{1}-m_{2}\right)\mathclose{}\rangle$
$\qquad+\frac{1}{2}\operatorname{Tr}_{X}\left[\mathopen{}\left(C_{2}+\gamma\mathbb{I}\right)\mathclose{}^{-1}\mathopen{}\left(C_{1}+\gamma\mathbb{I}\right)\mathclose{}-\mathbb{I}\right]$
$\qquad-\frac{1}{2}\log\operatorname{det}_{X}\left[\mathopen{}\left(C_{2}+\gamma\mathbb{I}\right)\mathclose{}^{-1}\mathopen{}\left(C_{1}+\gamma\mathbb{I}\right)\mathclose{}\right].$ (2.8)
Here $\operatorname{Tr}_{X}$ and $\operatorname{det}_{X}$ are the extended trace and extended Fredholm determinant [Quang, 2022]. For any $\gamma>0$, the regularized KL divergence is well-defined and finite (following Quang [2017, Proposition 1]), even if the Gaussian measures are singular [Quang, 2019]. It converges to the conventional KL divergence (if it is well-defined) for $\gamma\to 0$ (Quang, 2022, Theorem 6). Furthermore, if the Gaussian measures $\nu_{1}$ and $\nu_{2}$ are induced by GPs $\mathcal{GP}(\mu_{i},K_{i})$ for $i=1,2$, respectively, then $D^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\nu_{1}\,\|\,\nu_{2}\right)\mathclose{}$ is consistently estimated [Quang, 2022] by
$\hat{D}^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\nu_{1}\,\|\,\nu_{2}\right)\mathclose{}:=$ $\,\frac{1}{2}\mathopen{}\left({\bm{m}}_{1}-{\bm{m}}_{2}\right)\mathclose{}^{\top}({\bm{\Sigma}}^{(\gamma)}_{2})^{-1}\mathopen{}\left({\bm{m}}_{1}-{\bm{m}}_{2}\right)\mathclose{}$
$+$ $\,\frac{1}{2}\operatorname{Tr}\big{[}({\bm{\Sigma}}_{2}^{(\gamma)})^{-1}{\bm{\Sigma}}^{(\gamma)}_{1}-\mathbb{I}_{M}\big{]}$
-- $\,\frac{1}{2}\log\det\big{[}({\bm{\Sigma}}_{2}^{(\gamma)})^{-1}{\bm{\Sigma}}^{(\gamma)}_{1}\big{]}$ (2.9)
@@ 3 Generalized function-space VI with the regularized KL divergence
This section presents our proposed generalized function-space variational inference (GFSVI) method, which addresses the problem of the infinite KL divergence discussed in section 2.1, which we take for an indication that VI is too restrictive if one wants to use genuine function-space priors. We instead consider generalized variational inference [Knoblauch et al., 2022], which reinterprets the ELBO in Eq. 2.1 as a regularized expected log-likelihood and explores alternative divergences for the regularizer. Specifically, we propose to use the regularized KL divergence. This section builds heavily on tools introduced in section 2, which turn out to fit together perfectly: the pushforward of a Gaussian variational distribution in weight-space through the linearized neural network (Eq. 2.6) induces a GP variational posterior (Eq. 2.7) that admits a Gaussian measure on $L^{2}(\mathcal{X},\rho)$. Further, selecting a GP prior which has an associated Gaussian measure on $L^{2}(\mathcal{X},\rho)$ allows us to use the regularized KL divergence (Eq. 2.8). We present GFSVI in section 3.1 and compare it to prior work in section 3.2.
3.1 Generalized function-space VI
We present a well-defined objective for function-space inference, and a simple algorithm for its optimization.
Objective function.
We start from the ELBO in Eq. 2.3, where we use the Gaussian variational measure $\mathbb{Q}_{\phi}^{F}$ induced by the pushforward of a Gaussian variational distribution $q_{\phi}({\bm{w}})=\mathcal{N}\mathopen{}\left({\bm{w}}\,|\,{\bm{m}},{\bm{S}}\right)\mathclose{}$ along the linearized network $f_{L}$ (Eq. 2.6). The function-space prior may be any GP that has an associated Gaussian measure $\mathbb{P}^{F}$ on $L^{2}(\mathcal{X},\rho)$. We now replace the KL divergence in the ELBO with the regularized KL divergence $D_{\text{KL}}^{\gamma}$ (Eq. 2.8), which is well-defined and finite for any pair of Gaussian measures. For a likelihood function $p\mathopen{}\left(\mathcal{D}\,|\,{\bm{w}}\right)\mathclose{}=\prod_{i=1}^{N}p\mathopen{}\left(y_{i}\,|\,f_{L}({\bm{x}}_{i};{\bm{w}})\right)\mathclose{}$, we obtain
@@ 3.1 Generalized function-space VI
Estimation and optimization.
[…]
Technical details ($\gamma$ and $\rho$).
It turns out that increasing $\gamma$ reduces the influence of the prior on inference (see figure 20). At the same time, $\gamma$ acts as jitter that prevents numerical errors (see section 3.2). We recommend setting $\gamma$ large enough to avoid numerical errors but sufficiently small to strongly regularize the objective in Eq. 3.1 (see figure 18 in appendix) and setting $M$ to the largest value allowed by the computational budget. We found that the estimator $\hat{D}_{\text{KL}}^{\gamma}\big{(}\mathbb{Q}_{\phi}^{F}\,\big{|}\!\big{|}\,{\mathbb{P}^{F}}\big{)}$ converges quickly to a finite value (especially for smooth kernels, see figure 20 in appendix), and that GFSVI is robust to a wide range of values (we fixed $\gamma=10^{-10}$). The probability measure $\rho$ for $L^{2}(\mathcal{X},\rho)$ has to assign non-zero probability to any open set of $\mathcal{X}$ to regularize the BNN on all of its support. Following Rudner et al. [2022b], we draw measurement points from a uniform distribution over $\mathcal{X}$ when using tabular data and explore different configurations (samples from other data sets) for high-dimensional image data (see section C.4).
Figure 2: Inference on synthetic data (gray circles) using a Matérn-1/2 prior for function-space methods GFSVI and FVI. The proposed GFSVI provides the best approximation of the exact GP posterior.
@@
Linearized BNN $f_{L}$with measure $\mathbb{Q}_{\phi}^{F}$, GP prior $\mathcal{GP}(\mu,K)$with measure $\mathbb{P}^{F}$, measurement point distribution $\rho({\bm{x}})$, data $\mathcal{D}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}$, batch size $B$, $\gamma>0$. minibatch $({\bm{x}}_{\mathcal{B}},y_{\mathcal{B}})\sim\mathcal{D}$
1:Calculate ^$\hat{\ell}_{1}=\frac{N}{B}{\mathop{{}\mathbb{E}}}_{q_{\phi}({\bm{w}})}\mathopen{}\left[\log p\mathopen{}\left(y_{\mathcal{B}}\,|\,f_{L}({\bm{x}}_{\mathcal{B}},{\bm{w}})\right)\mathclose{}\right]\mathclose{}$;
2:Draw measurement points ${\bm{x}}=\{{\bm{x}}^{(i)}\}_{i=1}^{M}\stackrel{{\scriptstyle\!\!\text{i.i.d.}\!\!}}{{\sim}}\rho({\bm{x}})$;
3:Calculate ^$\hat{\ell}_{2}=\hat{D}_{\text{KL}}^{\gamma}\big{(}\mathbb{Q}_{\phi}^{F}\,\big{|}\!\big{|}\,{\mathbb{P}^{F}}\big{)}$ using ${\bm{x}}$ (Eq. 2.9);
4:Calculate ^$\hat{\mathcal{L}}(\phi)=\hat{\ell}_{1}-\hat{\ell}_{2}$
@@
[…]
@@ 3.2 Connections to prior work
TFSVI [Rudner et al., 2022b] and FVI [Sun et al., 2019] solve stability issues by introducing jitter/white noise, which has a similar effect as the regularization in Eq. 2.8. However, TFSVI introduces jitter only to overcome numerical issues and is fundamentally restricted to prior specification in weight space since its function-space prior is the pushforward of a weight-space prior. Conversely, FVI adds white noise to prevent the KL divergence (Eq. 2.4) to blow up as $M$ increases. However, FVI does not linearize the BNN, and hence does not have access to an explicit variational measure in function space. This severely complicates the estimation of (gradients of) the KL divergence in FVI, and the authors resort to implicit score function estimators, which make their method difficult to use in practice [Ma and Hernández-Lobato, 2021]. Our proposed GFSVI does not suffer from these difficulties as the variational posterior is an explicit Gaussian measure. This allows us to estimate the regularized KL divergence without sampling any noise or using implicit score function estimators.
@@ 4 Experiments
In this section, we evaluate our generalized function-space variational inference (GFSVI) method qualitatively on synthetic data and quantitatively on real-world data. GFSVI accurately captures structural properties specified by the GP prior, and that it performs competitively on regression, classification and out-of-distribution detection tasks. We also discuss the influence of the BNN’s inductive biases.
Baselines.
@@
We compare GFSVI to two weight-space inference methods: mean-field VI (MFVI) [Blundell et al., 2015] and linearized Laplace [Immer et al., 2021]; and to three function-space inference methods: FVI [Sun et al., 2019], TFVSI [Rudner et al., 2022b] and VIP [Ma et al., 2019] (TFSVI performs inference in function space but with the pushforward of a weight-space prior; VIP uses a BNN prior). All BNNs have the same architecture and fully-factorized Gaussian approximate posterior. We also include results for a sparse GP with a posterior mean parameterized by a neural network (GWI) [Wild et al., 2022b], and for a Gaussian Process (GP) [Williams and Rasmussen, 2006] (when the size of the dataset allows it), and for a sparse GP [Hensman et al., 2013] for regression tasks. We consider the GP, sparse GP and GWI as gold standards as they represent the exact (or near exact) posterior for models with GP priors.
Figure 3: Results for the ocean current modeling experiment. We report the norm of the mean velocity vectors and the squared errors. Unlike TFSVI, we find that GFSVI accurately captures ocean current dynamics.
Qualitative results on synthetic data.
@@
We consider a $1$-dimensional regression task where the values $y_{i}$ are sampled around $\sin(2\pi x_{i})$ (circles in Figures 1-8 and 11) and the two moons 2-dimensional binary classification task [Pedregosa et al., 2011] (see Figures 9 and 10). For regression, the green lines show functions sampled from the (approximate) posteriors, and the red lines are the inferred mean functions. For classification, the first and second row show the inferred mean probability of class 1 (blue dots) and its 2-standard deviations with respect to posterior samples. More details in Appendix C.1. We find that GFSVI captures the beliefs of the RBF and Matérn-1/2 GP priors better than BNN-baselines in the regression setting (see figures 5 and 2) as well as in classification (see figures 9 and 10), and shows greater uncertainty outside of the support of the data. figures 1 and 6 show that GFSVI notably adapts to varying prior assumptions (varying smoothness and length scale, respectively). In addition, figures 4 and 8 in the Appendix show that GFSVI provides strong regularization when the data generative process is noisy, and that it can be trained with fewer measurement points $M$ than FVI without significant degradation.
Inductive biases.
@@
figure 11 in the Appendix compares GFSVI to the exact GP-posterior across two different priors and three model architectures (details in Appendix C.1). We find that, with ReLU activations, small models are prone to underfitting for smooth priors (RBF), and to collapsing uncertainty for rough priors (Matérn-1/2). By contrast, with smooth activations (Tanh), smaller models suffice, and they are compatible with most standard GP priors (the results shown in figure 11 extend to RBF, Matérn, and Rational Quadratic in our experiments). We also analyzed how the number $M$ of measurement points affects performance. figures 7 and 17 in the appendix show that capturing the beliefs of rough GP priors and estimating $D_{\text{KL}}^{\gamma}$ with these priors requires larger $M$.
4.1 Quantitative results on real-world data